OpenStudy (anonymous):

How do you find the derivative of: (x^3 -3x^2 + 4) / x^2

OpenStudy (anonymous):

\[ (x^3 -3x^2 + 4) \div x^2\]

OpenStudy (anonymous):

I get, 1-8x^-3... so 1/-8x^3 Answer is x^3 - 8 / x^3

sam (.sam.):

You just split them \[\frac{ x^3 -3x^2 + 4 }{ x^2 } \\ \\ \\ \\ \frac{x^3}{x^2}-\frac{3x^2}{x^2}+\frac{4}{x^2} \\ \\ x-3+\frac{4}{x^2}\] Then use power rule

sam (.sam.):

\[\frac{d}{dx}x^n=nx^{n-1}\]

OpenStudy (anonymous):

That's exactly what I did.. \[1 - 8x^-3\]

sam (.sam.):

Your problem is on differentiating \[\frac{4}{x^2}\]

OpenStudy (anonymous):

that becomes (4) (-2x^-3)

sam (.sam.):

Careful, put it like this first \[\frac{4}{x^2} \\ \\ =4x^{-2}\] Then use power rule

sam (.sam.):

You get \(-8x^{-3}\)

OpenStudy (anonymous):

Yes

sam (.sam.):

Putting it all together \(1-8x^{-3}\)

OpenStudy (anonymous):

Yeeah, thats what I got 1-8x^-3

OpenStudy (anonymous):

Nvm... Got it.. 1 becomes x^3 / x^3